A car travels up a hill at a constant speed of 37 km/h and returns down the hill at a constant speed of 66 km/h. Calculate the average speed for the whole trip

Grade:11

2 Answers

Arun
4 years ago
 

By definition the average speed is the ration of the total traveled distance and the total traveled time. Let us introduce the total traveled distance of the car as L. Then the time of the travel up the hill is

t1 = L/74

The time of the travel down the hill is

t2 = L/132

The total traveled time is

 t = t1 + t2 = L/74 + L/132 = 0.0211 L

Then the average velocity is

 Vavg = L/t = L/0.0211 L = 47.393 km/hr

 
Vikas TU
4 years ago
Dear student 
Average speed Of thw whle trip will be 
= 2V1V2 / V1+ V2 
=> 2* 37 * 66 /103 
=> 47.47 m/sec 
Hope this clears 
Good Luck